# LakeVilla: Multi-Table Transactions for Lakehouses

### v1

@@ 5.3. Concurrent Workloads
OTFs and Concurrent Workloads. To evaluate the effects of concurrency and supporting multiple writers, we execute YCSB (Ren et al., 2024; Cooper et al., 2010) with 1, 8, and 64 concurrent clients and compare Spark [DL] with LV [R]. Spark was allowed up to 10 retries for a failed query. We evaluated LV [R] using single-query transactions (1Q Txn) and multi-query transactions that bundles all operations of a client (single Txn). We expect the throughput to increase for all workloads and systems, but also that the frequency of the conflicts will impact the behavior both through aborts by affecting the throughput. Figure 6 shows our results for YCSB A, B, C, D, and F. YCSB E was not tested because our current LV prototype does not support scans. Confirming our hypothesis, Spark experienced a slight increase in throughput, but with abort rates rising alongside the client count. Surprisingly, we also measured a small number of aborts for the read-only workload YCSB C, which we found to be an internal problem in Spark. In contrast, LV [R] had no aborts and outperformed Spark in throughput, particularly when bundling operations, which reduces version assignments per client. When we include the commit latency of LV [R] (single Txn) in our throughput calculations, we can also measure the impact for each workload type: For example, write-heavy workloads (A, F) experience a drop in throughput when we increase concurrency due to redo operations triggered at commit time. In contrast, executing LV [R] (1Q Txn) degenerates into normal Delta Lake with built-in retries, resulting in a throughput closer to Spark. As to be expected, bundling multiple operations into a transaction is advantageous for throughput as it minimizes the overhead of the version assignment phase. However, when overused, it may lead to penalties due to redo operations. In conclusion, we show that concurrency for writes is possible with LV [I], but users must be mindful of the suitable transaction sizes. The more writes a workload has, the less requests should be bundled into single transactions.
@@
Figure 7. YCSB (8 clients): Comparison of LV [I] and LV [R]
[…]
[YCSB (8 clients): Comparison of LV-I and LV-R]YCSB (8 clients): Comparison of LV [I] and LV [R]. LV [R] is faster for YCSB’s small reads and writes but more unstable than LV [I].
@@
LV[R] and LV[I]. Lastly, we use YCSB with varying workload types to compare the impact of LV[R] ’s markers and LV [I] ’s global version log on latency under concurrency. Again, we bundled all operations into one transaction per client and ran YCSB with eight clients. LV [I] ’s hidden version interval was set to two. We hypothesize that LV [R] will be more suitable for mixed workloads, whereas LV [I] will be better for read-heavy workloads. In Fig. 7, we report this experiment’s read and write latencies. They confirm our hypothesis in the following way: LV [R] provides better latency at the cost of more outliers. These are caused by marker acquisition failures, clean-up, and transaction retries. This is independent of the workload type, as reads and writes synchronize in the metadata layer. Conversely, LV [I] adds higher overhead overall but stabilizes the read operations, which can now only conflict with extremely outdated table versions (Fig. 7(b) and Fig. 7(c)). In write-heavy workloads (Fig. 7(a)), LV [I] caused more outliers due to failed validations requiring full transaction retries. In summary, LV [R] optimizes concurrent transaction synchronization for OTFs with faster performance but higher variability, while LV [I] stabilizes the reads under concurrency without introducing writer’s congestion.
5.4. Benefits and Trade-offs for Transactions
@@
Figure 8. Table Scaling of LV [CT]. The client in the table scaling experiment wrote one row to every opened table.
[…]
[Table Scaling of LV-CT. The client in the table scaling experiment wrote one row to every opened table.]Table Scaling of LV [CT]. The client in the table scaling experiment wrote one row to every opened table. LV [CT]’s latency requires at least 2 threads per table to compensate for its overhead. Afterwards, we see a linear increase.
@@
LV [CT] ’s trade-off. Next, we evaluate the strengths and performance trade-offs of LV [CT], which supports multiple tables. Hence, unlike LV [R], its overhead depends on the number of accessed tables per transaction. We scaled the number of tables, following the customer table definition of TPC-H. All transactions inserted a single row into each table to reduce operation overhead but ensure it triggers the LV [CT] ’s conflict detection across all tables. The LV clients use 64 threads, and our hypothesis is that a transaction’s commit latency will increase with the number of tables due to more metadata layers and markers being involved.
@@
Our results can be seen in Fig. 8. As expected, we can observe the overhead of LV [CT] ’s conflict detection, once we access more than one table. However, the transaction latency stays stable (at ca. 1s) until we run transactions accessing more than 32 tables. The parallelization in our LV prototype (and assigning 64 threads to our engine) causes this effect. Our prototype uses two threads per table for LV [CT]: one to execute the operations and the other to track dependencies to other tables. Hence, as soon as we run out of threads, some threads will need to manage two or more tables simultaneously, causing increased latency.
@@
Additionally, also the commit latency increases with each accessed table. The tracking of dependencies is the overhead of LV [CT], but this does not impact the overall latency for transactions with less than 32 tables. The expanded commit procedure of generating the log/snapshot entries and overwriting the markers for each accessed table separately only impacts this feature minimally.
@@
Figure 9. CAB-LV: Level Switching of read and write streams with their respective throughput on TPC-H data.
[…]
[CAB-LV: Level Switching of read and write streams with their respective throughput on TPC-H data.]CAB-LV: Level Switching of read and write streams with their respective throughput on TPC-H data. The best combinations for reads are LV [DL] and LV [I], while large writes benefit from LV [R] and LV [R, CT]. These recommendations hold under concurrency.
CAB-LV: Feature Combinations We now proceed with more complex workloads to investigate the effects on performance when we switch from one LV feature combination to another. To assess that, we adapted the Cloud Analytics Benchmark (CAB) (van Renen and Leis, 2023) for LHs and distinguish between read and write streams. Read streams execute the streams generated by the original CAB generator consisting of TPC-H queries. The write streams either insert up to 100 rows to a random table per query (80%) or issue update requests on a single table (20%). Our real-world observations on LH workloads inspired this distribution. We execute the experiment on TPC-H sf1, sf10, and sf100 and report the average latency of all metadata operations of all runs in Fig. 9. Each phase lasted 20 minutes until a client switches to a new feature combination.
CAB-LV: Sequential Runs. Our hypothesis for sequential runs of read and write streams follows our previous observations. Read will benefit from configurations without added congestion with writes, while writes can use marker-based combinations to minimize the aborts. In Fig. 9(a), we show the latency of sequential metadata operations. We observe the best performance for LV [DL] and LV [I], that avoid the overhead of marker placement. In contrast, the worst performance is observed when we enable all LV features. LV [CT] ’s conflict detection has an overhead when enabled on LV [R] or LV [I]. On its own, LV [CT] has a high latency, but benefits from other features that use markers (compare LV [R, CT] with LV [CT, I]).
@@
CAB-LV: Concurrent Runs. We tested our recommendations by executing reads and writes concurrently in our CAB scenario. Fig. 9(b) shows a plot of the latencies of the metadata operations. In this setting, LV [I] exhibits the lowest latency due to its global version assignment. In contrast, with LV [DL] the readers experience a higher overhead because version assignments are done for all tables individually. Concurrent writes using LV [R] are unaffected by concurrent reads with LV [I], and can thus benefit of the marker placement. Effectively, when compared to LV [DL], LV [R] moves the main synchronization to the beginning of a transaction with a minor impact on the performance and stability using its sublog structure. When adding LV [CT] to both operations, we can see more outliers for the writes, and slower reads. The readers now need to set markers within the table, making them compete with the writers. However, by using LV [CT] we enable all clients to detect concurrent transactions and react to them (by adapting or requesting a new version) before committing. In summary, we confirm our recommendations, but we also note that the effects were only noticeable because we isolated the metadata operations. If we also account for data loading, then the observable effects are minimal.
5.5. Analytical workloads
Figure 10. TPC-DS: Data loading times of Spark and LakeVilla
[…]
[TPC-DS: Data loading times of Spark and LakeVilla]TPC-DS: Data loading times of Spark and LakeVilla. LakeVilla can retrieve the required data faster than Spark. Its features show minimal overhead.
Analytics and LakeVilla. We also analyze the impact of LV on OLAP using three TPC-DS queries from (Jain et al., 2023) on scale factors 1000 and 3000. These queries stress the system as they rely heavily on OTF metadata design but also require selected data (Jain et al., 2023). Our experiment includes LV [R], LV [CT], LV [I], Spark’s cold, and Spark’s best-of-three (hot) performance. We measure how long each client actively interacts with the object store using MinIO’s tracing features. Additionally, we investigate data caching effects and their influence on OTFs by comparing Spark’s hot and cold runs. We expect LV’s changes in the metadata layer operations will minimally impact performance as data loading via the network is the main bottleneck. While LV, as a storage layer, can not fully orchestrate query execution on its own or fully replicate Spark’s task concurrency, our results can be seen as a lower bound for full systems. Further, we expect noticeable differences between Spark’s hot and cold runs due to avoided object store loads.
Our results in Fig. 10 confirm our hypothesis. Queries like Q67 primarily depend on data loading, while Q68 and Q90 spend more time on execution-driven or intermediate result caching by Spark. Overall, since all LV features show similar performance for all queries, we conclude that LV’s features either do not or only exhibit a minor impact on OLAP performance.
5.6. Compatibility: LakeVilla + Existing Systems
@@
Figure 11. Minimal additional requests needed for LakeVilla + Worst-case estimations for Delta Lake and Iceberg
[…]
[Minimal additional requests needed for LakeVilla + Worst-case estimations for Delta Lake and Iceberg.]Minimal additional requests needed for LakeVilla + Worst-case estimations for Delta Lake and Iceberg. Our estimations indicate that LakeVilla adds a minimal overhead in cost and possibly performance to Spark’s implementations.
@@
Figure 11 shows the characteristics of Spark’s Delta Lake and Iceberg with the added requests for each feature for their single-query transactions. LV’s markers and sublog add the most additional write requests. For reads (Fig. 11(a)), this impact becomes more apparent as for writes (Fig. 11(b)), where Spark already issues such requests. Thus, this further confirms our previous results of LV [R] being more suitable for write transactions. LV [CT] adds a fixed number of one write request and five reads per accessed table. Hence, like its latency, its costs depend on the number of accessed tables within a transaction. Lastly, LV [I] adds the least amount of additional requests, with its added write requests being skippable for read-only transactions in its WriteSerializable mode, integrating it well into read transactions.
6. Related Work
@@
Integrating and accessing data lakes. Previous research has focused on the operational efficiency of data lakes, like improving latency when accessing files by introducing strategies to encode relational data differently or improve file encodings (Kuschewski et al., 2023; Afroozeh and Boncz, 2023; Ghita et al., 2020). Further, Umbra focused on efficient parquet readers (Rey et al., 2023; Rey, 2024), transferring required sections of a parquet file and providing insights into the format structure. These approaches complement LV because Parquet files are commonly used for LH table data. Furthermore, cloud research has also explored elastic computing and additional structures to reduce cost and data lake traffic (Lee et al., 2023; Ritter, 2022; Vuppalapati et al., 2020; Winter et al., 2022; Durner et al., 2023; Perron et al., 2020; Müller et al., 2020; Kraft et al., 2023). These approaches follow the proposed architecture of external locks (Armbrust et al., 2020; Zaharia et al., 2021). In contrast, using a built-in approach, LV enables multiple features that previously required an auxiliary structure.
@@
Lakehouses. Multiple proposed designs enabled missing features (Chen et al., 2022; Madden et al., 2022; Levandoski et al., 2024) or specializations (Begoli et al., 2021; Hambardzumyan, 2023) in custom OTFs. They redesign the metadata layer, introduce file optimizations and orderings, or add auxiliary structures and provide built-in features like security or performance for neural network training. Unlike these, LV is an extension applicable to existing OTFs. Project Nessie implements git-like branching for Apache Iceberg to isolate different transactions (cf. (i) & (iii)) (licence, 2024). However, Project Nessie is a client catalog, while LV focuses on built-in solutions. Similarly, Photon explores design decisions for a vectorized query engine processing Delta Lake tables like the combined memory management or row- versus column-oriented processing (Behm et al., 2022). Photon’s insights can become relevant for our LV prototype in the future as both use Delta Lake.
@@
LakeVilla and External Solutions Multi-table transaction support is a heavily requested feature for LHs (Armbrust et al., 2020; Jain and Databricks, 2024; Foundation, 2024c, d), making it valuable to compare current advances to LV. The Delta Lake commit owner presumably is a hybrid solution, moving the commit decision to an external service and modifying Delta Lake to guarantee persistence (Jain and Databricks, 2024). In contrast, LV operates on the object store, making it a more independent solution. The possible advantages of the commit owners are better commit performance, tailored to Delta Lake, and cleaner metadata layers than LV’s markers. However, commit owners are presumably limited to a single vendor environment and not adaptive to other OTFs and transaction requirements. Additionally, the effects of an out-of-sync commit owner or the introduced ”backfilling” for compatibility are unknown and require further investigation (Jain and Databricks, 2024). We could see a combination of both used in practice: Commit Owners within a vendor’s environment for performance and LV for operations from the outside.
7. Conclusion

### v2

@@ 6.3 Concurrent Workloads
OTFs and Concurrent Workloads. To evaluate the effects of concurrency and supporting multiple writers, we execute YCSB conf/cloud/CooperSTRS10 ; YCSB-C with 1, 8, and 64 concurrent clients and compare Spark[DL] with LV[R] on S3 and MinIO. Spark was allowed up to 10 retries for a failed query. We evaluated LV[R] using single-query transactions (1Q Txn) and multi-query transactions that bundle all operations of a client (single Txn). We expect the throughput to increase for all workloads and systems, but also that the frequency of the conflicts will impact the behavior both through aborts by affecting the throughput. Figure 8 shows our results for YCSB A, B, C, D, and F. YCSB E was not tested because our current LV prototype does not support scans. Additionally, we only show the results for Spark[DL] on S3 as they represent both scenarios equally. Confirming our hypothesis, Spark experienced a slight increase in throughput, but with abort rates rising alongside the client count. Surprisingly, we also measured a small number of aborts for the read-only workload YCSB C, which we found to be an internal problem in Spark. In contrast, LV[R] had no aborts and outperformed Spark in throughput, particularly when bundling operations, which reduces version assignments per client. When we include the commit latency of LV[R] (single Txn) in our throughput calculations, we can also measure the impact for each workload type: For example, write-heavy workloads (A, F) experience a drop in throughput when we increase concurrency due to redo operations triggered at commit time. In contrast, executing LV[R] (1Q Txn) degenerates into normal Delta Lake with built-in retries, resulting in a throughput closer to Spark. As to be expected, bundling multiple operations into a transaction is advantageous for throughput as it minimizes the overhead of the version assignment phase. However, when overused, it may lead to penalties due to redo operations. We observed these results for both S3 and MinIO. In the S3 case, higher network latency decreased the overall throughput, mostly penalizing bundling operations into a single transaction. In conclusion, independent of the used object store, we show that concurrency for writes is possible with LV[I], but users must be mindful of the suitable transaction sizes. The more writes a workload has, the fewer requests should be bundled into single transactions.
@@
Figure 9: YCSB (8 clients, S3): Comparison of LV[I] and LV[R] (matplotlib)
@@
LV[R] and LV[I]. Lastly, we use YCSB with varying workload types to compare the impact of LV[R] ’s markers and LV[I] ’s global version log on latency under concurrency. Again, we bundled all operations into one transaction per client and ran YCSB with eight clients using S3. LV[I] ’s hidden version interval was set to two. We hypothesize that LV[R] will be more suitable for mixed workloads, whereas LV[I] will be better for read-heavy workloads. In Figure 9, we report this experiment’s read and write latencies. They confirm our hypothesis in the following way: LV[R] provides better write latency at the cost of more outliers. These are caused by marker acquisition failures, clean-up, and transaction retries. LV[R] adds the same overhead to reads and writes as both synchronize in the metadata layer. Conversely, LV[I] shows lower read overhead and stabilized performance, because reads can now only conflict with extremely outdated table versions. For writes (Figures 9(a) and 9(b)), LV[I] shows a larger overhead than LV[R] due to its commit procedure and failed validations requiring full retries. In summary, LV[R] optimizes concurrent writes with faster performance but higher variability, while LV[I] is ideal for stable reads under concurrency without introducing writer’s congestion.
6.4 Benefits and Trade-offs for Transactions
@@
Figure 10: Table Scaling of LV[CT] on MinIO. The client wrote one row to every opened table (matplotlib)
@@
LV[CT] ’s trade-off. Next, we evaluate the strengths and performance trade-offs of LV[CT], which supports multiple tables. Unlike LV[R], LV[CT]’s overhead depends on the number of accessed tables per transaction. We scaled the number of tables on MinIO, following the customer table definition of TPC-H. All transactions inserted a single row into each table to reduce operation overhead but ensure it triggers LV[CT]’s conflict detection across all tables. The LV clients use 64 threads, and we hypothesize that a transaction’s commit latency will increase with the number of tables due to more metadata layers and markers being involved.
@@
Our results can be seen in Figure 10. As expected, we can observe the overhead of LV[CT] ’s conflict detection once we access more than one table. However, the transaction latency stays stable (at ca. 1s) until we run transactions accessing more than 32 tables. The parallelization in our LV prototype (and assigning 64 threads to our engine) causes this effect. Our prototype uses two threads per table for LV[CT]: one to execute the operations and the other to track dependencies to other tables. Hence, as soon as we run out of threads, some threads will need to manage two or more tables simultaneously, causing increased latency.
@@
Additionally, the commit latency increases with each accessed table. The tracking of dependencies is the overhead of LV[CT], but this does not impact the overall latency for transactions with less than 32 tables. The expanded commit procedure of generating the log/snapshot entries and overwriting the markers for each accessed table separately impacts this feature minimally.
@@
Figure 11: TPC-H (S3, sf10) throughput of different LV features of CAB read and write streams journals/pvldb/RenenL23 (matplotlib)
CAB-LV: Feature Combinations We now proceed with more complex workloads to investigate the effects on performance when we switch from one LV feature combination to another. To assess that, we adapted the Cloud Analytics Benchmark (CAB) journals/pvldb/RenenL23 for LHs and distinguished between read and write streams. Read streams execute the streams generated by the original CAB generator, consisting of TPC-H queries. The write streams either insert up to 100 rows into a random table per query (80%) or issue update requests on a single table (20%). Our real-world observations on LH workloads inspired this distribution. We execute the experiment on TPC-H sf10 on S3 and report their latency in Figure 11. Each phase lasted 20 minutes until a client switches to a new feature combination.
CAB-LV: Sequential Runs. Our hypothesis for sequential runs of read and write streams follows our previous observations. Reads will benefit from configurations without added congestion with writes, while writes can use marker-based combinations to minimize aborts. In Figure 11(a), we show the latency of sequential executions. We observe the best performance for LV[DL] and LV[I], which avoid the overhead of marker placement. In contrast, the worst performance is observed when we enable all LV features. LV[CT] ’s conflict detection has an overhead when enabled on LV[R] or LV[I]. On its own, LV[CT] has a high latency, but benefits from other features that use markers (compare LV[R, CT] with LV[CT, I]).
@@
CAB-LV: Concurrent Runs. We tested our recommendations by executing reads and writes concurrently in our CAB scenario. Figure 11(b) shows a plot of the latencies of the metadata operations. In this setting, LV[I] exhibits the lowest latency due to its global version assignment. In contrast, with LV[DL] the readers experience a higher overhead because version assignments are done for all tables individually. Concurrent writes using LV[R] are unaffected by concurrent reads with LV[I], and can thus benefit from the marker placement. Effectively, when compared to LV[DL], LV[R] moves the main synchronization to the beginning of a transaction with a minor impact on the performance and stability using its sublog structure. When adding LV[CT] to both operations, we can see more outliers for the writes and slower reads. The readers now need to set markers within the table, making them compete with the writers. However, by using LV[CT], we enable all clients to detect concurrent transactions and react to them (by adapting or requesting a new version) before committing. In summary, we confirm our recommendations, but we also note that the effects were only noticeable because we used a smaller scale factor (sf10). For larger amounts of data, we must account more for data loading, making observable effects minimal.
6.5 Transactional and Mixed Workloads
Table 4: TPC-C with 1 warehouse and sequential and concurrent clients[successful txns/min] on S3
Transaction type sequential concurrent
single-op LV[DL] (4.22)
single-table LV[R] (3.42)
multi-table
LV[CT] (3.49) (2.79)
LV[I] (3.67) (2.81)
= no abots     = aborts occurred     = impossible
Analytical Workloads. Further, we analyzed pure OLTP and mixed workloads. To confirm the capabilities of our individual LV features, we tested TPC-C tpcc using one warehouse to increase the occurrence of transactional conflicts. We hypothesize that LV[CT] and LV[I] can execute full TPC-C transactions while showing aborts. In contrast, while working fine when run sequentially, LV[DL] and LV[R] cannot execute TPC-C under concurrency. Table 4 shows our results for sequential and concurrent (16 clients) runs, confirming our hypothesis. All configurations were able to execute successfully sequentially. However, LV[DL] and LV[R] could not do a single TPC-C transaction without any consistency violations. LV[CT]’s and LV[I]’s mechanisms successfully detected consistency violations and triggered aborts for erroneous transactions. Although LV[CT] and LV[I] are able to execute full OLTP workloads, they also showcase that LHs generally are not ideal for such scenarios due to their low throughput.
(a) Throughput
(b) First three repetitions of SU and DM
Figure 12: LSTBench (scale factor 100): W0 and W3 with LV[R, CT] (writes) and LV[I] (reads) on S3 (matplotlib)
Figure 13: TPC-DS (scale factor 1000): Data loading times of Spark and LakeVilla on MinIO. All entries marked with ’*’ do not use data partitioning due to technical limitations. (matplotlib)
@@ 6.5 Transactional and Mixed Workloads
Figure 14: MinIO statistics: Minimal additional requests needed for LV in Delta Lake and Iceberg (matplotlib)
Mixed Workloads. After one extreme, we analyzed typical LH scenarios using LSTBench’s W0 and W3 2024lstbench on TPC-DS scale factor 100 with a selection of 50 random TPC-DS queries (out of Q1-Q10, Q67, Q68, Q90). Following our previous results, we use LV[R, CT] for Data Maintenance (DM), LV[CT, I] for Optimize (O), and LV[I] for SingleUser (SU) tasks. As W0 and W3 focus on concurrent analytics, LV should be able to provide the desired guarantees with minimal impact on performance over a long time. Figure 12 shows our results when using S3. As theorized, the throughput for respective phases, shown in Figure 12(a), benefits from concurrent SU phases (4xSU in W0) but is impacted by updates introduced by DM phases. O phases compactify the metadata layers for reads in SU, but can not compensate for added data. The compactification effect benefits SU when the metadata is large after multiple updates (compare the first and second O in W3).
Further, when inspecting the latency of the first three occurrences of SU and DM in W0 and W3 in Figure 12(b), we see two types: metadata-intensive (below 1s SUs and DMs) and data-intensive transactions (SU). Metadata-intensive operations primarily interact with the LH metadata layer and stay stable throughout runs. In contrast, data-intensive transactions load extensive data and are impacted by an increasing data size. We can observe this effect with the shift of the largest/second largest bar for SU to the right, the later the W0 run is executed. While LV can keep the Metadata layer at a reasonable size (see O phases), it can not compensate for increased data sizes. We theorize that an integration of LV in analytical systems would improve stability among phases with a built-in cache, as is common for systems interacting with LHs.
6.6 Analytical workloads
Analytics and LakeVilla. Lastly, we analyze the impact of LV on pure OLAP using three TPC-DS queries from conf/cidr/0001KPDSZ23 on scale factor 1000. These queries rely heavily on OTF metadata design, require selected data, and show the most significant difference across the tested systems conf/cidr/0001KPDSZ23 . Our experiment includes LV[R], LV[CT], LV[I], Spark’s and DuckLake’s cold, and Spark’s and DuckLake’s best-of-three (hot) performances. Using the tracing features of MinIO, we measure how long each client interacts with it. Additionally, we investigate data caching effects and their influence on OTFs by comparing hot and cold runs. We expect LV’s changes in the metadata layer operations will minimally impact performance, as data loading via the network is the main bottleneck. While LV, as a storage layer, can not fully orchestrate query execution or fully replicate DuckLake’s and Spark’s task concurrency, our LV results present a lower bound for complete systems. Further, we expect noticeable differences between DuckLake’s and Spark’s hot and cold runs due to avoided object store loads. As DuckLake provides its metadata layer within DuckDB, we expect it to be faster than all alternatives, as it must only load the real data from MinIO.
Our results in Figure 13 mostly confirm our hypothesis. Queries like Q67 primarily depend on data loading, while Q68 and Q90 spend more time on execution-driven or intermediate result caching by Spark and DuckLake. However, DuckLake’s cold performance was slower than expected due to the system not partitioning the data on MinIO. It’s hot runs for Q68 and Q90 (where it only partially loads the data from MinIO) validate this problem, as they are on par with Spark and LV. However, for Q67 that requires a huge amount of data, DuckLake’s cache is unable to compensate for its simplistic data partitioning on the object store, affecting even its hot runs. Overall, since all LV features show similar performance for all queries, we conclude that LV’s features either do not or only exhibit a minor impact on OLAP.
6.7 Compatibility: LakeVilla + Existing Systems
@@
Figure 14 shows the characteristics of Spark’s Delta Lake and Iceberg with the added requests for each feature for their single-query transactions. LV’s markers and sublog add the most additional write requests. For reads (Figure 14(a)), this impact becomes more apparent as for writes (Figure 14(b)), where Spark already issues such requests. Thus, this further confirms our previous results of LV[R] being more suitable for write transactions. LV[CT] adds a fixed number of one write request and five reads per accessed table. Hence, like its latency, its costs depend on the number of accessed tables within a transaction. Lastly, LV[I] adds the least amount of additional requests, with its added write requests being skippable for read-only transactions in its WriteSerializable mode, integrating it well into read transactions.
7 Related Work
@@
Integrating and accessing data lakes. Previous research has focused on the operational efficiency of data lakes, like improving latency when accessing files by introducing strategies to encode relational data differently or improve file encodings journals/pvldb/AfroozehB23 ; journals/pacmmod/KuschewskiSAL23 ; conf/cidr/GhitaTB20 . Further, Umbra focused on efficient parquet readers conf/edbt/Rey24 ; conf/btw/ReyFN23 , transferring required sections of a parquet file and providing insights into the format structure. These approaches complement LV because Parquet files are commonly used for LH table data. Furthermore, cloud research has also explored elastic computing and additional structures to reduce cost and data lake traffic journals/pvldb/DurnerL023 ; journals/pvldb/Kraft0ZBSYZ23 ; conf/damon/LeeWADBS0KR23 ; conf/sigmod/0002MA20 ; conf/sigmod/PerronFDM20 ; journals/is/Ritter22 ; conf/nsdi/VuppalapatiMATM20 ; journals/pvldb/WinterGNK22 . These approaches follow the proposed architecture of external locks journals/pvldb/ArmbrustDPXZ0YM20 ; conf/cidr/Zaharia0XA21 . In contrast, using a built-in approach for LHs, LV enables multiple features that previously required an auxiliary structure.
@@
Lakehouses. Multiple proposed designs enabled missing features conf/cbd/ChenSLLJ22 ; conf/icde/DeyFR15 ; conf/sigmod/LevandoskiCDDEH24 ; conf/cidr/MaddenDKSCMT22 or specializations conf/bigdataconf/BegoliGK21 ; conf/cidr/Hambardzumyan23 in custom OTFs or other setups. They redesign the metadata layer, introduce file optimizations and orderings, or add auxiliary structures and provide built-in features like security or performance for neural network training. Unlike these, LV is an extension applicable to existing OTFs. Project Nessie implements git-like branching for Apache Iceberg to isolate different transactions (cf. (i) & (iii)) project_nessie_web . However, Project Nessie is a client catalog, while LV focuses on built-in solutions. Similarly, Photon explores design decisions for a vectorized query engine processing Delta Lake tables like the combined memory management or row- versus column-oriented processing conf/sigmod/BehmPAACDGHJKLL22 . Photon’s insights can become relevant for our LV prototype in the future as both use Delta Lake.
@@
LV and External Solutions Multi-table transaction support is a heavily requested feature for LHs journals/pvldb/ArmbrustDPXZ0YM20 ; delta4_multi_table ; iceberg_multi_table ; iceberg_nessie_doc ; ducklake , making it valuable to compare current advances to LV. The Delta Lake commit owner is a hybrid solution, moving the commit decision to an external service and modifying Delta Lake to guarantee persistence delta4_multi_table . In contrast, LV operates on the object store, making it a more independent solution. The possible advantages of the commit owners are better commit performance, tailored to Delta Lake, and cleaner metadata layers than LV’s markers. However, commit owners are presumably limited to a single vendor environment and not adaptive to other OTFs and transaction requirements. Additionally, the effects of an out-of-sync commit owner or its ”backfilling” for compatibility are unknown and require further investigation delta4_multi_table .
DuckLake moves the Lakehouse metadata layer into DuckDB, using the object store purely for table data ducklake . This approach reuses DuckDB features and makes LHs more lightweight ducklake . However, we found limitations of this approach while preparing our experiments: First, DuckLake currently does not allow the use of the same LH by multiple sessions and processes inherited by DuckDB’s design duckdb_concurrency , only allowing sequential transactions. Further, it does not synchronize data changes outside DuckDB or upholds any structure on the object store, making DuckDB a single bottleneck and the data unusable for any other system with access to the object store. Lastly, DuckLake uses no data partitioning on the object store, affecting its performance (see TPC-DS measurements in Section 6). While DuckLake excels in scalability, it ignores some crucial design aspects seen in previous OTF designs, focused on by LV.
Future Combinations. Hence, we could see a combination of the three used in practice: DuckLake or Commit Owners within a vendor’s environment for performance and LV for operations from the outside.
8 Conclusion